OpenStudy (anonymous):

5. Heinz boasts that he can predict the degree of a polynomial function just by looking at the end behavior. Can Heinz do this? Explain.

OpenStudy (ash2326):

@Katt7 what do you mean by end behavior?

OpenStudy (anonymous):

end behavior is like the odd or even degree in the polynomial function @ash2326

OpenStudy (kmeis002):

Well, you have began to partiall answer your question. You have said that the end behavior can signify odd or even highest degrees, such as \[f(x) = x^{3} \] approaches -infinity from the left and positive from the right. However, I would ask you to graph \[f(x) = x^3 + 3x \] and \[f(x) = x^5 - 5x^2 + x\] Just by looking at both end behaviours, would you be able to tell what degree the function has?

OpenStudy (agent0smith):

No, you can't predict the degree - only whether the highest power of x is odd or even.